From a given cost function
TC =60+0.5q+0.3q×q
find out fixed cost, average variable cost and marginal cost
1
Expert's answer
2015-04-13T09:19:20-0400
TC =60 + 0.5q + 0.3q×q
Fixed cost FC = 60, as it doesn't change at different quantities.
Average variable cost AVC = VC/q = (TC - FC)/q = (0.5q + 0.3q×q)/q = 0.5 + 0.3q
Marginal cost MC = change in TC/change in q = TC' = 0.5 + 0.6q